OpenStudy (anonymous):

domain of y=tan(8x) in mathematical notation?

OpenStudy (barrycarter):

Can you do it in non-math notation?

OpenStudy (anonymous):

I don't know how to do it either way.

OpenStudy (barrycarter):

OK, what numbers *can't* you put in for 8x?

OpenStudy (anonymous):

I know I'm supposed to use pi and that's all I really know.

OpenStudy (barrycarter):

Where does tan() have asymptotes?

OpenStudy (anonymous):

pi/2?

OpenStudy (barrycarter):

Yes! Every pi/2. So where does tan(8x) [note the 8x] have asymptotes?

OpenStudy (anonymous):

Oh! so the domain should be like, can't = pi/2, and so on?

OpenStudy (anonymous):

I think...

OpenStudy (barrycarter):

The domain of tan(x) excludes all pi/2 plus all multiplies of 2pi, and also -pi/2 plus all multiples of 2pi. But this is tan(8x)

OpenStudy (anonymous):

so it should just be can't = 8pi/2?

OpenStudy (barrycarter):

Afraid not. It's where 8x = pi/2 + 2*pi*n or 8x=-pi/2+2*pi*n

OpenStudy (anonymous):

I have no clue

OpenStudy (barrycarter):

well, try dividing what I typed for 8x by 8 to get x=(something) or x=(something else)